# Diffusion approximation of the stationary distribution of a two-level single server queue

### v7

@@ 6.1 Process limit
We now tentatively assume that there exists finite $\widetilde{p}_{1,\ell_{0}}$ such that
$\widetilde{p}_{1,\ell_{0}}=\lim_{r\downarrow 0}\frac{p^{(r)}_{1,\ell^{(r)}_{0}}}{r}.$ (A.26)
The existence of this $\widetilde{p}_{1,\ell_{0}}$ will be verified at the final step (see (A.31)). Applying these asymptotic properties, it follows from (A.2) and (A.20) with help of (A.26) that
$\lim_{r\downarrow 0}g_{2}^{(r)}(0)$ $=\lim_{r\downarrow 0}\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}+\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{1-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{0}}$
$=\frac{c_{1}\mu+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{2,\ell^{(r)}_{0}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{1,\ell^{(r)}_{0}}=0,$ (A.27)
$\lim_{r\downarrow 0}g^{(r)}_{3}(0)$ $=\lim_{r\downarrow 0}\frac{1}{1-\rho^{(r)}_{2}}p^{(r)}_{3,\ell^{(r)}_{0}+1}=\frac{c_{2}}{b_{2}}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{3,\ell^{(r)}_{0}+1}$
$=\beta_{2}^{-1}\gamma_{22}^{-1}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{2,\ell^{(r)}_{0}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\beta_{2}^{-1}\widetilde{p}_{1,\ell_{0}}.$ (A.28)
@@
$g^{(r)}_{1}(0)$ $=\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(0)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(0)\right)\frac{(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad+\left(f^{(r)}_{1}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}.$ (A.29)
@@
$\lim_{r\downarrow 0}rf^{(r)}_{1}(0)=\lim_{r\downarrow 0}r\ell^{(r)}_{0}=\ell_{0},$
$\lim_{r\downarrow 0}rf^{(r)}_{\rho_{1}}(0)=\lim_{r\downarrow 0}r\frac{1-(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}}{1-\rho^{(r)}_{1}}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{0}}),$
$\lim_{r\downarrow 0}f^{(r)}_{\gamma_{12}^{-1}}(0)(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}=\lim_{r\downarrow 0}\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{1-\gamma^{(r)}_{12}}\gamma^{(r)}_{12}=\frac{c_{1}\mu}{\lambda_{2}}=\rho_{21}^{-1}.$
@@
$\lim_{r\downarrow 0}\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{0}})\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{1,0}$
$\quad=\beta_{1}^{-1}\left(e^{\beta_{1}\ell_{0}}-1\right)\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\widetilde{p}_{1,\ell_{0}},$
$\lim_{r\downarrow 0}\left(f^{(r)}_{1}(0)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(0)\right)\frac{(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{0}}=0,$
$\lim_{r\downarrow 0}\left(f^{(r)}_{1}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}=0,$
@@
$\lim_{r\downarrow 0}\left(f^{(r)}_{\gamma^{-1}_{12}}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad=\lim_{r\downarrow 0}f^{(r)}_{\gamma^{-1}_{12}}(0)(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}p^{(r)}_{2,\ell^{(r)}_{0}}=0.$
@@
$\lim_{r\downarrow 0}g^{(r)}_{1}(0)=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\beta_{1}^{-1}(e^{\beta_{1}\ell_{0}}-1)\widetilde{p}_{1,\ell_{0}}.$ (A.30)
@@
$1=\lim_{r\downarrow 0}g^{(r)}(0)=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\left(\beta_{2}^{-1}+\beta_{1}^{-1}(e^{\beta_{1}\ell_{0}}-1)\right)\widetilde{p}_{1,\ell_{0}}.$ (A.31)
This proves that $\widetilde{p}_{1,\ell_{0}}$ indeed exists and is finite, and given by
$\widetilde{p}_{1,\ell_{0}}=\frac{\lambda_{2}\beta_{1}\beta_{2}}{(\lambda_{1}+\lambda_{2})(\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1))}.$ (A.32)
We are now in a final step to find $\widetilde{g}(\theta)\equiv\lim_{r\downarrow 0}g^{(r)}(\theta)$ for $\theta<0$. This can be obtained from (A.2), (A.2) and (A.20) in the following way. From (A.20) and Lemma A.1,
$\lim_{r\downarrow 0}g^{(r)}_{3}(\theta)$ $=\lim_{r\downarrow 0}\frac{1}{1-e^{r\theta}\rho^{(r)}_{2}}e^{(\ell^{(r)}_{0}+1)r\theta}p^{(r)}_{3,\ell^{(r)}_{0}+1}$
$=\lim_{r\downarrow 0}\frac{r}{1-(1+r\theta+o(r\theta))(1-r\beta_{2}+o(r))}e^{\ell^{(r)}_{0}\theta}\frac{1}{r}p^{(r)}_{3,\ell^{(r)}_{0}+1}$
$=\frac{e^{\ell_{0}\theta}}{\beta_{2}-\theta}\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{1,\ell^{(r)}_{0}}$
$=\frac{e^{\ell_{0}\theta}\beta_{2}}{\beta_{2}-\theta}\frac{\beta_{1}}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)},$
@@
$\lim_{r\downarrow 0}g^{(r)}_{1}(\theta)$ $=\lim_{r\downarrow 0}\left(f^{(r)}_{\rho_{1}}(\theta)+(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}=\lim_{r\downarrow 0}\frac{1-(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}+1}}{1-e^{r\theta}\rho^{(r)}_{1}}p^{(r)}_{1,0}$
$=\frac{1-e^{\ell_{0}(\theta-\beta_{1})}}{\beta_{1}-\theta}\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}e^{\beta_{1}\ell_{0}}\widetilde{p}_{1,\ell_{0}}$
$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\ell_{0}\theta}}{\beta_{1}-\theta}\frac{\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}$
Thus, we have, for $\theta\leq 0$ satisfying $\theta\not=\beta_{1}$,
@@
$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\ell_{0}\theta}}{\beta_{1}-\theta}\frac{\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}+\frac{e^{\ell^{(r)}_{0}\theta}\beta_{2}}{\beta_{2}-\theta}\frac{\beta_{1}}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{0}}-1)}.$
@@
$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}]$ $=\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,1})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)\leq\ell^{(r)}_{0}]$
$\quad+\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,2})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)>\ell^{(r)}_{0}],$
@@
$e^{r\theta\ell^{(r)}_{0}}=e^{\theta(\ell_{0}+o(r))}=e^{\theta\ell_{0}}(1+\theta o(r))+o(\theta o(r))=e^{\theta\ell_{0}}(1+o(\theta r)),$
$e^{r\theta(\ell^{(r)}_{0}+1)}=e^{\theta r\ell^{(r)}_{0}}e^{r\theta}=e^{\theta\ell_{0}}(1+o(\theta r))(1+\theta r+o((\theta r)^{2}))=e^{\theta\ell_{0}}(1+\theta r+o(\theta r)).$
@@
$e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}=1-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}+\frac{1}{2}(\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-})^{2}+O\big{(}|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|^{3}\max_{r|\theta|<a}e^{|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|}\big{)}$
@@
A.5 Proofs of Lemmas 5.1 and 5.2
We first prove Lemma 5.1, namely, that (5.3) implies (5.4). To prove this, we use the BAR (4.1) for $f(z,x_{1},x_{2})=(x_{1}\wedge 1/r)(x_{2}\wedge 1/r)1(z=\ell^{(r)}_{0})$, then we have
@@
Hence, (5.4) follows from (A.5) because $\mathbb{E}[\widehat{T}^{(r)}_{s}]$ converges to $\mu^{-1}$ as $r\downarrow 0$. Thus, Lemma 5.1 is proved.
We next prove Lemma 5.2. By Schwartz’s inequality,
$\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0)=\ell^{(r)}_{0}]\big{\}}^{2}=\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3/2}(\widehat{R}^{(r)}_{e})^{1/2}1^{2}(L^{(r)}(0)=\ell^{(r)}_{0})]\big{\}}^{2}$
$\quad\leq\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{0})]\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})].$
Since $\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{0})]$ is uniformly bounded in $r$ by Corollary 4.4, (5.4) implies (5.5). This complete the proof of Lemma 5.2.

### v8

@@ 6.1 Process limit
We now tentatively assume that there exists finite $\widetilde{p}_{1,\ell_{1}}$ such that
$\widetilde{p}_{1,\ell_{1}}=\lim_{r\downarrow 0}\frac{p^{(r)}_{1,\ell^{(r)}_{1}}}{r}.$ (A.27)
The existence of this $\widetilde{p}_{1,\ell_{1}}$ will be verified at the final step (see (A.32)). Applying these asymptotic properties, it follows from (A.3) and (A.20) with help of (A.27) that
$\lim_{r\downarrow 0}g_{2}^{(r)}(0)$ $=\lim_{r\downarrow 0}\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}+\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}}{1-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{1}}$
$=\frac{c_{1}\mu+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{2,\ell^{(r)}_{1}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}p^{(r)}_{1,\ell^{(r)}_{1}}=0,$ (A.28)
$\lim_{r\downarrow 0}g^{(r)}_{3}(0)$ $=\lim_{r\downarrow 0}\frac{1}{1-\rho^{(r)}_{2}}p^{(r)}_{3,\ell^{(r)}_{1}+1}=\frac{c_{2}}{b_{2}}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{3,\ell^{(r)}_{1}+1}$
$=\beta_{2}^{-1}\gamma_{22}^{-1}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{2,\ell^{(r)}_{1}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\beta_{2}^{-1}\widetilde{p}_{1,\ell_{1}}.$ (A.29)
@@
$g^{(r)}_{1}(0)$ $=\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{1}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(0)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(0)\right)\frac{(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
$\quad+\left(f^{(r)}_{1}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}.$ (A.30)
@@
$\lim_{r\downarrow 0}rf^{(r)}_{1}(0)=\lim_{r\downarrow 0}r\ell^{(r)}_{1}=\ell_{1},$
$\lim_{r\downarrow 0}rf^{(r)}_{\rho_{1}}(0)=\lim_{r\downarrow 0}r\frac{1-(\rho^{(r)}_{1})^{\ell^{(r)}_{1}}}{1-\rho^{(r)}_{1}}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{1}}),$
$\lim_{r\downarrow 0}f^{(r)}_{\gamma_{12}^{-1}}(0)(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}=\lim_{r\downarrow 0}\frac{1-(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}}{1-\gamma^{(r)}_{12}}\gamma^{(r)}_{12}=\frac{c_{1}\mu}{\lambda_{2}}=\rho_{21}^{-1}.$
@@
$\lim_{r\downarrow 0}\left(f^{(r)}_{\rho_{1}}(0)+(\rho^{(r)}_{1})^{\ell^{(r)}_{1}}\right)p^{(r)}_{1,0}=\beta_{1}^{-1}(1-e^{-\beta_{1}\ell_{1}})\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{1,0}$
$\quad=\beta_{1}^{-1}\left(e^{\beta_{1}\ell_{1}}-1\right)\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\widetilde{p}_{1,\ell_{1}},$
$\lim_{r\downarrow 0}\left(f^{(r)}_{1}(0)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(0)\right)\frac{(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}=0,$
$\lim_{r\downarrow 0}\left(f^{(r)}_{1}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}=0,$
@@
$\lim_{r\downarrow 0}\left(f^{(r)}_{\gamma^{-1}_{12}}(0)-f^{(r)}_{\rho_{1}}(0)\right)\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
$\quad=\lim_{r\downarrow 0}f^{(r)}_{\gamma^{-1}_{12}}(0)(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}p^{(r)}_{2,\ell^{(r)}_{1}}=0.$
@@
$\lim_{r\downarrow 0}g^{(r)}_{1}(0)=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\beta_{1}^{-1}(e^{\beta_{1}\ell_{1}}-1)\widetilde{p}_{1,\ell_{1}}.$ (A.31)
@@
$1=\lim_{r\downarrow 0}g^{(r)}(0)=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\left(\beta_{2}^{-1}+\beta_{1}^{-1}(e^{\beta_{1}\ell_{1}}-1)\right)\widetilde{p}_{1,\ell_{1}}.$ (A.32)
This proves that $\widetilde{p}_{1,\ell_{1}}$ indeed exists and is finite, and given by
$\widetilde{p}_{1,\ell_{1}}=\frac{\lambda_{2}\beta_{1}\beta_{2}}{(\lambda_{1}+\lambda_{2})(\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1))}.$ (A.33)
We are now in a final step to find $\widetilde{g}(\theta)\equiv\lim_{r\downarrow 0}g^{(r)}(\theta)$ for $\theta<0$. This can be obtained in the following way. From (A.20) and Lemma A.1,
$\lim_{r\downarrow 0}g^{(r)}_{3}(\theta)$ $=\lim_{r\downarrow 0}\frac{1}{1-e^{r\theta}\rho^{(r)}_{2}}e^{(\ell^{(r)}_{1}+1)r\theta}p^{(r)}_{3,\ell^{(r)}_{1}+1}$
$=\lim_{r\downarrow 0}\frac{r}{1-(1+r\theta+o(r\theta))(1-r\beta_{2}+o(r))}e^{\ell^{(r)}_{1}\theta}\frac{1}{r}p^{(r)}_{3,\ell^{(r)}_{1}+1}$
$=\frac{e^{\ell_{1}\theta}}{\beta_{2}-\theta}\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}\lim_{r\downarrow 0}\frac{1}{r}p^{(r)}_{1,\ell^{(r)}_{1}}$
$=\frac{e^{\ell_{1}\theta}\beta_{2}}{\beta_{2}-\theta}\frac{\beta_{1}}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1)},$
@@
$\lim_{r\downarrow 0}g^{(r)}_{1}(\theta)$ $=\lim_{r\downarrow 0}\left(f^{(r)}_{\rho_{1}}(\theta)+(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{1}}\right)p^{(r)}_{1,0}=\lim_{r\downarrow 0}\frac{1-(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{1}+1}}{1-e^{r\theta}\rho^{(r)}_{1}}p^{(r)}_{1,0}$
$=\frac{1-e^{\ell_{1}(\theta-\beta_{1})}}{\beta_{1}-\theta}\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}e^{\beta_{1}\ell_{1}}\widetilde{p}_{1,\ell_{1}}$
$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\ell_{1}\theta}}{\beta_{1}-\theta}\frac{\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}.$
Further, $\lim_{r\downarrow 0}g^{(r)}_{2}(\theta)=0$ by (A.3). Hence, we have, for $\theta\leq 0$ satisfying $\theta\not=\beta_{1}$,
@@
$=\frac{\beta_{1}}{e^{\beta_{1}\ell_{1}}-1}\frac{e^{\beta_{1}\ell_{1}}-e^{\ell_{1}\theta}}{\beta_{1}-\theta}\frac{\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}+\frac{e^{\ell^{(r)}_{1}\theta}\beta_{2}}{\beta_{2}-\theta}\frac{\beta_{1}}{\beta_{1}+\beta_{2}(e^{\beta_{1}\ell_{1}}-1)}.$
@@
$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}]$ $=\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,1})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)\leq\ell^{(r)}_{1}]$
$\quad+\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,2})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)>\ell^{(r)}_{1}],$
@@
$e^{r\theta\ell^{(r)}_{1}}=e^{\theta(\ell_{1}+o(r))}=e^{\theta\ell_{1}}(1+\theta o(r))+o(\theta o(r))=e^{\theta\ell_{1}}(1+o(\theta r)),$
$e^{r\theta(\ell^{(r)}_{1}+1)}=e^{\theta r\ell^{(r)}_{1}}e^{r\theta}=e^{\theta\ell_{1}}(1+o(\theta r))(1+\theta r+o((\theta r)^{2}))=e^{\theta\ell_{1}}(1+\theta r+o(\theta r)).$
@@
$e^{-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}}$
$\quad=1-\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}+\frac{1}{2}(\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-})^{2}+O\big{(}|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|^{3}\max_{r|\theta|<a}e^{|\zeta^{(r)}(r\theta)\widehat{R}^{(r)}_{s-}|}\big{)}$
@@
A.6 Proof of Lemma 5.1
We first assume $\lambda_{1}\not=\lambda_{2}$, and prove (5.3) for $i=1$. Consider the BAR (4.3) for $\theta_{2}=\theta_{1}=\theta\leq 0$. From (4.2), we have
$D^{(r)}(\theta,\theta)=\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{1}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (A.38)
zz Applying (4.38), (4.8) and Corollary 4.4 to this equation, we have
$D^{(r)}(\theta,\theta)$ $=\alpha^{(r)}_{e}e^{\theta\ell_{1}}(1+r\theta+o(r\theta))\mathbb{E}^{(r)}_{s}\big{[}(r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\widehat{R}^{(r)}_{e}$
$\quad-2^{-1}r^{2}\theta^{2}\{(\lambda^{(r)}_{2})^{2}-(\lambda^{(r)}_{1})^{2}\}(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{1})\big{]}+o((r\theta)^{2}),$
$=\alpha^{(r)}_{e}e^{\theta\ell_{1}}\big{\{}r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})\big{]}+(r\theta)^{2}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})\big{\}}$
$\quad+o((r\theta)^{2}).$
Since ${\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2}$ is uniformly bounded in $r\in(0,1]$, this implies that
$D^{(r)}(\theta,\theta)$ $=r\theta\alpha^{(r)}_{e}e^{\theta\ell_{1}}(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})\big{]}+O((r\theta)^{2}).$ (A.39)
$\quad=c_{1}\left(\mu r\theta-\mu^{3}\sigma_{s}^{2}r^{2}\theta^{2}/2+o((\theta r)^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+D^{(r)}(\theta,\theta).$ (A.40)
Then, picking up the terms of order $r\theta$, we have
$c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{\theta\ell_{1}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})]=O(r\theta).$
$(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})]=O(r).$ (A.41)
Since $\lambda_{1}\not=\lambda_{2}$, (A.41) implies (5.3) for $i=1$. To prove (5.3) for $i=2$, we derive the following formula by Schwartz’s inequality.
$\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{2}1(L^{(r)}(0)=\ell^{(r)}_{1}]\big{\}}^{2}=\big{\{}\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3/2}(\widehat{R}^{(r)}_{e})^{1/2}1^{2}(L^{(r)}(0)=\ell^{(r)}_{1})]\big{\}}^{2}$
$\quad\leq\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{1})]\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{1})].$ (A.42)
Since $\mathbb{E}^{(r)}_{s}[(\widehat{R}^{(r)}_{e})^{3}1(L^{(r)}(0)=\ell^{(r)}_{1})]$ is uniformly bounded in $r$ by Corollary 4.4, (5.3) for $i=1$ implies that for $i=2$. Thus, (5.3) is proved.
We next prove that (5.4) implies (5.3). For this proof, we use the BAR (4.1) for $f(z,x_{1},x_{2})=(x_{1}\wedge 1/r)(x_{2}\wedge 1/r)1(z=\ell^{(r)}_{0})$, then we have
@@
Hence, we have (5.3) for $i=1$. It for $i=2$ follows from (A.6) as we have argued.
A.7 Proof of Lemma 5.2
In the proof of Proposition 5.2, under the assumption $\lambda_{1}\not=\lambda_{2}$, (5.2) is obtained applying (5.3) to (5.2). Obviously, if (5.3) holds, then this proof is also valid without any extra assumption including $\lambda_{1}\not=\lambda_{2}$. Hence, the lemma is proved.